OpenStudy (anonymous):

Find another name for the given number, but with the denominator indicated. use multiplication by 1. 5/6=?/36

hero (hero):

1.5/6 = n/36 This equation has the form a/b = c/d. Anytime you have an equation of this form, you can cross multiply to get ad = bc In this case, we have 1.5/6 = n/36. We can cross multiply to get 36(1.5) = 6n Solving for n we get: 54 = 6n 54/6 =n 9 = n

OpenStudy (anonymous):

Thank you!